OpenStudy (anonymous):

a^5/a^3 A. a^2 B. a^1.666 C. a^15

OpenStudy (anonymous):

answer is A. becaus a^5/a^2=a^(5-2)=a^3

OpenStudy (anonymous):

um.. The right answer is B. because you are dividing 5 over 3.